### v2

# GAR: Generative Adversarial Reinforcement Learning for Formal Theorem Proving

Abstract
Solving math problems through verifiable languages such as Lean has significantly impacted both the mathematics and computer science communities. Current state-of-the-art models are often trained with expensive online Reinforcement Learning (RL) or expert iteration. However, these approaches rely on fixed problem sets, which causes inefficient training and limits the model to tackle complex problems. To overcome these limitations, we propose GAR: Generative Adversarial Reinforcement learning, a comprehensive RL training framework that jointly trains the problem composer and solver in an adversarial loop. GAR introduces an implicit curriculum learning mechanism, which aligns task difficulty with the prover’s evolving capability. It thereby improves the training efficiency and enables stronger performance of proving advanced theorems. Experiments show that with GAR training, Goedel-Prover-V2-8B and DeepSeek-Prover-V2-7B achieve an average relative improvement in pass@32 of 4.20% on MiniF2F-Test benchmark, while DeepSeek-Prover-V2’s pass@32 on ProofNet-Test increases from 22.58% to 25.81%. Beyond formal proving, GAR establishes a general RL paradigm for co-evolution of problem generation and solving under verifiable environments. The training code for this paper is open-sourced in https://github.com/RickySkywalker/GAR-Official
1 Introduction
The capability to perform formal mathematical reasoning has long been regarded as both a hallmark of human intelligence and a key objective of machine learning (Newell & Simon, 1956). The ability is typically assessed through rigorous mathematical derivations (Yang et al., 2024b). With the emergence of Large Language Models (LLMs), developing accurate and reliable reasoning has become an active area of research. Recent progress in ZERO RL training (Guo et al., 2025) has further advanced reasoning systems by introducing Long Chain-of-Thought (CoT) thinking models that have self-reflection and self-correction capability.
However, the inherent ambiguity of Natural Language (NL) makes it challenging to verify intermediate reasoning steps. This problem is more severe in advanced mathematics, where the task is to prove theorems rather than give a numerical or formulaic answer. The increasing complexity of modern math compounds this difficulty, as illustrated by the prolonged verification of Fermat’s Last Theorem (Wang et al., 2024). To address this issue, researchers have grounded reasoning in formal logical systems, enabling automatic verification through Formal Language (FL). Based on this idea, some researchers model the reasoning process formally with dependent type languages like Lean (De Moura et al., 2015; Moura & Ullrich, 2021) and Coq (Coq, 1996). Other uses higher-order logic to build language like Isabelle (Paulson, 1994) and HOL (Harrison, 2009). All the above languages make every intermediate step of math reasoning verifiable.
Nevertheless, mastering FLs requires considerable expertise, particularly in dependent-type systems like Lean, where proofs often demand complex type matching and the use of functions with limited data (Wang et al., 2025c). Thus, many works have sought to leverage advances in LLMs to solve FL problems and train specialized FL provers (Polu et al., 2022; Jiang et al., 2022; Xin et al., 2024a; Wang et al., 2024; Lin et al., 2025a; Dong & Ma, 2025b). The verifiability of FLs also motivates large-scale synthesis of new statements, which enables extensive expert iteration (Polu et al., 2022; Xin et al., 2025) or Reinforcement Learning (RL) (Ren et al., 2025; Wang et al., 2025a; Lin et al., 2025b) to further enhance perver’s performance.
However, most prior approaches in RL or expert iteration rely on fixed theorem statement sets and optimize only the prover. The statement dataset is unchanged during the process. It also leads to wasted computation on trivial or unsolvable tasks. In expert iteration, datasets often require repeated annotation. On the other hand, advanced RL methods suffer from the absence of an adaptive difficulty level of statements during rollouts, which restricts concentrated exploration and limits progress on complex theorems. More broadly, only a few works discuss a systematic way to align problem difficulty with the prover’s growing skill.
To address this limitation, we propose GAR: Generative Adversarial Reinforcement Learning, a comprehensive RL training framework as presented in Figure 1. GAR jointly optimizes the prover and the problem composer, termed as the statement fuser, through adversarial training. This process establishes an implicit curriculum learning that dynamically adjusts statement difficulty to match the prover’s development. Each iteration of the GAR framework consists of two stages: the generation and the adversarial RL stage. In the generation stage, the statement fuser synthesizes harder statements from existing solvable ones, and the prover attempts to solve them. In the RL training stage, the fuser is rewarded for producing difficult yet solvable statements, while the prover is rewarded for correctly proving medium and high-difficulty problems. This adversarial setup ensures that statement complexity evolves alongside the prover’s capabilities, resulting in more efficient and effective training.
We summarize our contributions as follows: (1) We introduce GAR, a comprehensive RL training framework that establishes implicit curriculum learning, improving the prover’s reasoning ability while preventing wasted effort on trivial or unsolvable statements. (2) We present Statement Fusion technique, which enables the creation of novel formal statements beyond direct formalization of NL problems, producing theorem statements better aligned with model capability. (3) We demonstrate the effectiveness of the GAR through extensive experiments by training two base models using GAR. We achieve an average of 4.20% relative improvement on MiniF2F-Test (Zheng et al., 2021) and raise DeepSeek-Prover-8B’s pass@32 on the more challenging ProofNet-Test (Azerbayev et al., 2023) from 22.58% to 25.81%. It highlights the effectiveness and generality of GAR.
Furthermore, the GAR contributes a general RL paradigm for the co-evolution of problem generation and solving in verifiable environments, which offers new directions for adversarial co-training in other reasoning-intensive domains. To facilitate the advancement of the field, we will open-source the training and inference code of GAR in the near future.
2 Methodology
In this section, we introduce the GAR framework in detail. The prover’s task is to produce Lean4 proofs from NL-FL statement pairs. Our framework trains the prover to improve at the task by adapting theorem difficulty to the prover’s evolving capabilities through adversarial training of both the prover and the statement fuser. Each iteration of GAR consists of two stages: the Generation stage (Section 2.1) and the Adversarial Reinforcement Learning stage (Section 2.2). We first describe the behavior of each stage in a single iteration in detail, then present the complete procedure in Section 2.3.
2.1 Generation Stage
The generation stage of GAR consists of two processes, namely statement fusion, which produces new statements, and proof writing, which generates and evaluates candidate proofs.
2.1.1 Statement Fusion
This process generates harder statements from existing ones. It aims to mitigate the mismatch between the fixed datasets and the prover’s evolving ability. This process first samples two NL statements from the base dataset, formally: , where is a repository consists of 793,243 NL statements collect from Lean-Workbook (Wu et al., 2024b) and Numina-Math (LI et al., 2024). The sampled pair is then passed to the statement fuser trained in the previous iteration of GAR. The fuser is instructed to synthesize a more challenging statement by combining the key ideas of the two inputs, namely:
where denotes the current iteration index, is the generated NL statement, and is the fuser from the prior step (initialized as base model when ).
In particular, we chose thinking models like Qwen3 (Team, 2025) or DeepSeek-R1 (Guo et al., 2025) as the base model for the statement fuser because of its outstanding performance. However, direct use of their native Long CoT capability often results in overthinking and degraded quality of the generated statement. To mitigate this, we reinitialize the thinking process by passing the default thinking stage and restarting it with a dedicated indicator token named analysis. This trick yields more focused and task-specific reasoning. The full prompt for the statement fusion is shown in Figure 2. Subsequently, the generated NL statement is passed to the autoformalizer, which converts the NL statement into Lean statement . The formalized statement is then passed to Lean4 for compilation check. This step filters out statements with grammatical errors.
In the statement fusion stage, they deliberately separate the fusion of NL statements from the formalization of FL statements. Two key insights guide this design. First, prior studies (Wang et al., 2024; Xin et al., 2024b) have shown that NL statements play a crucial role in supporting formal reasoning. Secondly, current 8B-scale general LLMs exhibit only a limited understanding of FL. This causes them to fail to capture essential relations, which hinders the generation of more robust statements. Consequently, the fuser produces statements that have a high likelihood of failing the compilation check if we directly fuse formal statements, as seen in Dong & Ma (2025b). Thus, GAR firstly fuses the statement in NL and formalizes it to FL, ensuring the fused statements are adaptively more challenging than the base problems to fit prover performance evolution.
2.1.2 Proof Writing
Utilizing the NL-FL statement pair generated by the fusion process, the prover from the last iteration generates candidate proofs (with in our implementation), denoted as , specifically:
When , it is the base prover model like deepSeek-Prover-V2 (Ren et al., 2025) or Goedel-Prover-V2 (Lin et al., 2025b).
The set of candidate proofs is then passed to the Lean4 verifier to check correctness. Based on the verification results, we compute the empirical pass rate of this statement for reward assignment as , where is the number of proofs that pass verification.
2.2 Adversarial Reinforcement Learning
This section outlines the training process of GAR, where the prover and statement fuser are optimized in competition, deriving adversarial learning and mutual improvement.
2.2.1 Statement Fuser Training
The statement fuser is trained to synthesize problems that are slightly beyond the prover’s current capability, thereby creating an implicit curriculum. To achieve this, we assign rewards to the generated NL statements that encourage the production of more challenging but solvable problems. We adopt a variant of Group Relative Policy Optimization (GRPO) (Shao et al., 2024) for this optimization. Specifically, for each step , the fuser generates statements (with in our implementation). Formally, generated statement set is: , and base statement set is represented as: . The statement fuser is then updated by maximizing the following objective.
where denotes the policy model with parameter . It represents the statement fuser here. is the parameter for base statement fuser, is the parameter for the fuser in the previous step, and are hyper-parameters, and is the advantage, computed from the reward set by:
where is the prover’s empirical pass rate on , and is the statement modification rate, which indicates the portion of proofs that the prover tries to modify its statement. The reward is set to if the prover fails to solve the problem, which indicates the statement is too difficult or unsolvable.
We introduce the term as a soft statement modification penalty to balance the risk between reward hacking and the need to preserve model capability. Because of Long CoT training, current expert provers often acquire strong self-correction capability. While valuable, this ability can lead the model to change formal statements during proof writing. This may lead to serious reward hacking if unconstrained. Conversely, a strict ban on such modifications would suppress self-correction and reduce proofreading accuracy. Our soft penalty discourages excessive statement change without too much harm to the result.
In summary, the training design for the statement fuser rewards the model to lower the prover’s pass rate by composing more challenging statements.
2.2.2 Prover Training
To achieve adverasrial training, the prover is optimized to maxmize the pass rate on generated statements. We employ a variant of the GRPO algorithm for such training. Specifically, for each theorem statement , we update the prover model by maximizing the following objective function:
where is the prover as policy model, is the parameter of prover, and is the advantage of proof , computed from the reward set by:
where is a binary indicator of whether a statement modification occurred in proof . Similar to the training of the statement fuser, modifications are penalized but not strictly prohibited. Furthermore, to ensure the prover is trained on high-quality data, we exclude statements with an empirical pass rate of (unsolvable) or above (too easy) following experience in Wang et al. (2025a); Dong & Ma (2025b). It make sure the prover is only trained with hard and medium-level problems. This optimization scheme drives the prover to compete with the statement fuser by continually enhancing its proof-generation capability.
2.3 Summary
The GAR is an iterative framework (Wang et al., 2023) that loops the generation stage and the RL stage details above. To provide a clear overview, we present the complete framework in the form of pseudo-code as follows:
When trained in multiple iterations, the GAR establishes an implicit curriculum by aligning statement difficulty with the prover’s evolving capability. The statement fuser is trained to reduce the pass rate by generating more challenging statements, while the prover is optimized to increase it by producing more valid proofs. Together, their adversarial interaction drives progressive improvement.
3 Experiments
We conduct comprehensive experiments on the MiniF2F-Test (Zheng et al., 2021) and ProofNet-Test (Azerbayev et al., 2023) benchmarks to assess the performance of the GAR framework in formal proof writing. Specifically, we show in Section 3.3 that the models trained with GAR achieve better empirical results, confirm in Section 3.4 that adversarial training induces an implicit curriculum by generating progressively harder statements, and report ablation study results in Section 3.5. Due to space limitations, we have included the efficiency study and case study in Appendix D.
3.1 Experiment Setup
3.1.1 Dataset and Task
We measure the Lean4 reasoning capability of the GAR trained model by MiniF2F-Test (Zheng et al., 2021) and ProfNet-Test (Azerbayev et al., 2023) benchmarks. They are challenging benchmarks and adopted in nearly all major studies in the field (Xin et al., 2024b; Lin et al., 2024; Wang et al., 2024; Wu et al., 2024a; Polu et al., 2022; Dong & Ma, 2025a; Lin et al., 2025b; Wang et al., 2025c).
| Method | Sample budget | MiniF2F-Test | ProofNet-Test |
| Lean-STaR (Lin et al., 2024) | 46.31% | - | |
| InternLM-2.5-StepProver (Wu et al., 2024a) | 50.70% | 18.80% | |
| DeepSeek-Prover-V1.5-RL (Xin et al., 2024b) | 128 | 50.00% | 18.20% |
| STP-Lean (Dong & Ma, 2025b) | 128 | 56.15% | 19.50% |
| MA-LoT (Wang et al., 2025c) | 32 | 61.07% | 15.47% |
| Kimina-Prover-Distill-7B (Wang et al., 2025a) | 32 | 63.10% | - |
| DeepSeek-Prover-V2-7B (Ren et al., 2025) | 32 | 70.49% | 22.58% |
| Geodel-Prover-V2-8B (Lin et al., 2025a) | 32 | 77.87% | - |
| Our models | |||
| GAR on Deepseek-Prover-V2 | 32 | 74.18% | 25.81% |
| GAR on Goedel-Prover-V2 | 32 | 80.33% | - |
| Step idx | Base Model | GAR model |
| 0 | 29.16% | 29.16% |
| 1 | 16.50% | 23.71% |
| 2 | 11.58% | 20.53% |
| 3 | 7.61% | 20.08% |
| 4 | 7.69% | 21.79% |
| Step idx | w/o Stat. Penalty | Full GAR |
| 0 | 42.94% | 42.94% |
| 1 | 48.18% | 48.72% |
| 2 | 60.42% | 30.50% |
| 3 | 71.82% | 39.65% |
| 4 | 74.11% | 33.63% |
The MiniF2F-Test benchmarks comprise 244 Lean4 statements, spanning from high school competition problems to elementary undergraduate-level theorem proofs. Specifically, MiniF2F-Test comprises problems formalized from the MATH dataset (Hendrycks et al., 2021), high school competitions such as AMC, AIME, and IMO, as well as self-crafted problems. ProofNet-Test consists of 186 theorems formalized from standard undergraduate textbooks on advanced topics such as real and complex analysis, linear algebra, abstract algebra, and topology. In our setting, we train the LLM with GAR to generate Lean4 proofs from the NL-FL statement pair. To avoid overloading the model, all the imports and namespaces are manually configured.
3.1.2 Baselines
To evaluate the effectiveness of GAR, we compare it against strong open-source baselines, including Lean-STaR (Lin et al., 2024), InternLM-2.5-StepProver (Wu et al., 2024a), Kimina-Prover-Preview-Distill-7B (Wang et al., 2025a), DeepSeek-Prover-V1.5-RL (Xin et al., 2024b), STP-Lean (Dong & Ma, 2025a), MA-LoT (Wang et al., 2025c), Goedel-Prover-V2 (Lin et al., 2025b), and DeepSeek-Prover-V2 (Ren et al., 2025). For baseline models that require Long CoT reasoning (Kimina-Prover, MA-LoT, DeepSeek-Prover-V2, and Goedel-Prover-V2), we restrict the reasoning length to 16,384 tokens to conserve computational resources.111Because of this restriction, our reported results for DeepSeek-Prover-V2 and Goedel-Prover-V2 differ from those in the original papers, where the evaluations used a context length of 40,960.
3.2 Implementation Details
In the generation stage, we construct a repository of 793,243 NL statements from Numina-Math (LI et al., 2024) and Lean-Workbook (Ying et al., 2024) datasets. For statement fusion, we employ Qwen3-8B (Team, 2025) as the base model for the statement fuser because of its skip-thinking capability. We apply Kimina-Autoformalizer-7B (Wang et al., 2025a) as the autoformalizer. For proof generation, we use DeepSeek-Prover-V2-7B (Ren et al., 2025) and Goedel-Prover-V2 (Lin et al., 2025b) as base models for provers. We sample 1,024 theorems per step and generate 16 proofs per theorem following (Wang et al., 2025a). We also restrict the sequence length to 16,384 tokens for both models. The GRPO hyperparameters are set with a learning rate of 2E-6, , and . We perform three iterations of GAR training on Goedel-Prover-V2 and five iterations on DeepSeek-Prover-V2, which costs around 140 H100 hours for each training. On average, DeepSeek-Prover-V2 costs fewer hours per iteration due to average shorter thinking. In verification, any proof involving statement modification or relying on the “sorry” or “admit” tactic is counted as incorrect to ensure fairness.
3.3 Results
Table 1 demonstrates the empirical results of applying GAR to train base prover models. For Goedel-Prover-V2, MiniF2F-Test pass@32 raises to 80.33%, indicating a 3.16% relative gain. For DeepSeek-Prover-V2, the MiniF2F-Test score improves to 74.18%, corresponding to a 5.23% increase. On the more challenging ProofNet-Test benchmark, which targets advanced mathematics topics, GAR enhances DeepSeek-Prover-V2’s pass@32 rate from 22.58% to 25.81%.222We do not report Goedel-Prover-V2’s accuracy on ProofNet due to the absence of the reference results in Lin et al. (2025b).
Compared with other baselines, both the base model and the GAR-trained models achieve consistent and substantial gains, highlighting that GAR can contribute to models that already have outstanding performance. Compared to base models, the improvements suggest that the adversarial method proposed in GAR enables provers to tackle increasingly difficult problems by establishing an implicit curriculum. Such a curriculum can progressively adapt problem difficulty to the model’s capability, allowing the prover to explore deeper reasoning strategies and solve more advanced theorems. This behavior becomes more significant as the theorems get harder.
3.4 Statement Fuser Study
To show that GAR produces increasingly more difficult statements across iterations and thereby establishes an implicit curriculum learning, we examine the difficulty of problems generated by the statement fuser. In each iteration, we randomly sample 50 generated statements and compute their average proof correctness rate using both the base prover and GAR-trained prover at that iteration. Details of the metric we used are provided in Appendix B.2. The result of this experiment is reported in Table 3.
From the results, we can observe that the base model exhibits a consistent performance degradation, with accuracy falling from 29.16% at the first iteration’s data to 7.69% by the fifth. In contrast, the GAR-trained model shows only a minor initial decline due to the statement matching but stabilizes at around 21% across later iterations. These findings confirm that GAR progressively generates harder statements, as evidenced by the base model’s decline. The maintained trained model’s performance demonstrates the adversarial training strength of the prover’s performance over time.
3.5 Ablation Studies
3.5.1 Effect of statement modification penalty
We evaluate the impact of the statement modification penalty by monitoring statement modification rates across by the provers during training rounds. Details of this metric are provided in Appendix B.3. We trained a variant of GAR on Goedel-Prover-V2-8B without the penalty in both the statement fuser and prover, and compared it with the full GAR.
| Method | MiniF2F-Test |
| Base model | 77.87% |
| GRPO trained | 77.46% |
| GAR trained | 80.33% |
Table 3 shows that without the modification penalty, the prover exploits its self-correction ability by simplifying statements. Such behavior worsens as the training progresses; by the fourth step, 74% of statements were modified at least once, which is a clear signal of reward hacking. In contrast, with the penalty enabled, the modification rate remains below 40% throughout. This study confirms the effectiveness of our statement penalty in preventing reward hacking.
3.5.2 Compare to Direct RL training
This experiment tests whether jointly evolving the prover and statement fuser is more effective than training the prover alone using existing data. We conducted three additional GRPO iterations on Goedel-Prover-V2-8B using the same sampling budget and formalized NL base dataset as GAR. The results are shown in Table 4. The GAR-tained model outperforms the variant trained with traditional GRPO. We conclude that such an improvement to GAR progressively raises problem difficulty, enabling the prover to handle more complex statements. In contrast, further RL training on static dataset degrades performance, as the base model is already heavily RL trained. These findings further support the generality of GAR, demonstrating is ability to enhance model’s performance while standard RL no longer provides benefits.
4 Related Work
4.1 LLM for Formal Theorem Proving
Recently, applying LLMs to support formal theorem proving has become a prominent research direction. Training approaches can be broadly divided into two categories: model trained solely with Supervised Fine-Tuning (SFT) and those combining SFT with Reinforcement Learning (RL). Early SFT-based provers include Expert Iteration (Polu & Sutskever, 2020), Re-Prover (Yang et al., 2024c), TheoremLlama (Wang et al., 2024), DeepSeek-Prover-V1 (Xin et al., 2024a), InternLM-2.5-StepProver (Wu et al., 2024a), MA-LoT (Wang et al., 2025c), and Goedel-Prover-V1 (Lin et al., 2025a). These models typically require multiple rounds of large-scale annotation with existing formal solvers, which demands a significant amount of computational resources and limits exploration. To further advance the prover, researchers began incorporating advanced RL techniques. For instance, DeepSeek-Prover-V1.5 (Xin et al., 2024b) employs DPO. After the ”ZERO” RL technique that enables Long CoT thinking developed by Guo et al. (2025), provers like Kimina-Prover (Wang et al., 2025a), DeepSeek-Prover-V2 (Ren et al., 2025), and Goedel-Prover-V2 (Lin et al., 2025b) enable models to produce better formal proof after thinking. However, existing RL approaches still rely on a fixed collection of statements, preventing statements from adapting to the prover’s evolving skill. In contrast, GAR jointly trains a statement fuser and a prover, ensuring that generated theorems remain suitably challenging as the model improves.
4.2 RL methods for LLM
RL has been central to the development of reasoning LLMs (Xiong et al., 2025). Early efforts such as DeepSeek-Math (Shao et al., 2024) and Qwen-2.5-Math (Yang et al., 2024a) applied reward modeling with GRPO algorithm to enhance exploration. The release of OpenAI-O1 (OpenAI, 2024) and DeepSeek-R1 (Guo et al., 2025) demonstrated that verifier-based rewards in RL can enable complex reasoning strategies, including backward search and self-correction. Such techniques have been widely adopted in systems like Qwen3 (Team, 2025) and Kimi-K2 (Team et al., 2025). Nevertheless, nearly all prior work uses the verification signal only to optimize the problem solver, leaving the problem composer untrained. On the other hand, GAR simultaneously improves both statement fuser and the prover, creating an implicit curriculum that adapts task difficulty and allows the model to acquire more complex reasoning skills step-by-step.
5 Conclusion
This paper presents GAR, a comprehensive training framework for formal theorem proving. GAR aims to resolve the inefficient and suboptimal performance caused by traditional expert iteration and online RL in prover training. GAR achieves more efficient training by using statement fusion to formulate statements in the RL process to avoid the prover from annotating problems beyond its capability range. After GAR training, provers can solve more advanced problems through the implicit curriculum learning that enables the model to explore further. Furthermore, GAR can improve the performance of models that have been through heavy RL training by letting the model explore more difficult statements during the training process. Experiments of applying GAR to DeepSeek-Prover-V2 and Goedel-Prover-V2 yield an average relative improvement of 4.20% on the MiniF2F-Test dataset and improve the DeepSeek-Prover-V2’s performance on ProofNet-Test from 22.58% to 25.81%. Beyond theorem proving, GAR offers a general RL paradigm of co-evolution of the problem generation and solving under a verifiable environment. It provides a foundation for adversarial co-training in other reasoning-intensive domains.
Ethics Statement
After carefully reviewing the ethical regulations of the conference, to the best of our knowledge, this work does not present any foreseeable ethical concerns. This research focuses exclusively on formal theorem proving using publicly available mathematical datasets, without involving human subjects, private data, or sensitive content. As far as we are concerned, no negative societal or ethical impacts are anticipated for the contribution of this work. We only use LLMs to polish the writing style and fix grammatical errors in the paper.
Reproducibility Statement
We have made efforts to ensure that our work is reproducible. The detailed description of the GAR framework, including pseudocode, data source, reward definitions, and training objectives, is presented in Section 2. The experimental setup, including benchmarks, baselines, base model choice, and hyperparameters, is presented in Section 3.1 and Section 3.2. We also plan to further ensure the reproducibility by open-sourcing the code in the near future.
Appendix A Acknowledgment
This material is based upon work supported partially by NSF under Grant No. 2416897, Grant No. 2505932, and by ORN under Grant No. N000142512318. This research used both Delta (NSF award OAC 2005572) and DeltaAI (NSF award OAC 2320345) advanced computing systems, and computing resources provided by NAIRR Pilot NAIRR250157.
Appendix B Details of the metrics
For completeness, this section provides the metric definitions that were omitted due to space limitations in Section 3.
B.1 Pass Rate
Pass@ is a widely used metric for evaluating formal theorem provers (Polu et al., 2022; Jiang et al., 2022; Wang et al., 2024; Xin et al., 2024a; Dong & Ma, 2025b; Wang et al., 2025b; Lin et al., 2025b; Wu et al., 2024a; Wang et al., 2025a; Li et al., 2025; Yao et al., 2025; Zhang et al., 2026). For each theorem statement, the model generates candidate proofs, which are further checked by the Lean4 verifier. If at least one of the candidates is correct, the theorem is counted as a proved theorem. The pass@ of the LLM is the fraction of theorems in the dataset with at least one correct proof generated by the prover.
B.2 Average proof correctness rate
This metric measures the relative difficulty of a set of statements for a given prover, which is used in Section 3.4). Let be a dataset of size . For each theorem , we sample 16 proofs from the prover and record the number of proofs that pass the verification, which is . The average proof correctness rate is calculated by
A high value indicates the dataset is easier for the given prover.
Note that in this experiment, to demonstrate the significance of GAR better, we train two extra rounds of Goedel-Prover-V2-8B model.
B.3 Statement modification rate
Since current advanced provers have Long Chain-of-thought (CoT) capability, which enables self-reflection, backtracking, and self-correction. In the Long CoT process, a prover may alter the original statement into a simplified variant in its reasoning trace and ultimately produce a proof for the modified version. To quantify this behavior, we define statement modification rate as the portion of theorems in a dataset that have at least one statement modification across 16 generated proofs. This metric helps assess how often the prover attempts to modify the problem and whether such a behavior may be too significant to cause severe reward hacking.
Appendix C Direct comparison with Related Works
This section aims to provide a more direct comparison between GAR and other works that also try to build a dynamic dataset during training, namely Goedel-Prover-V2 (Lin et al., 2025b) and STP (Dong & Ma, 2025b).
Compared to Goedel-Prover-V2:
The statement generation in Goedel-V2 relies on a frozen large model to synthesize data for SFT. Crucially, during the RL phase, their statement set remains static. This lack of dynamic updates of statements based on the prover’s evolving capabilities may lead to suboptimal performance and efficiency as the prover outpaces the fixed problem set. In contrast, GAR continuously updates the statement fuser via adversarial RL, ensuring the generated statements remain progressively challenging and aligned with the prover’s current skill level during the RL phase, leading to better empirical performance.
Compared to STP:
While STP trains a conjecture model based on the prover’s feedback, it relies on the expert iteration framework based on offline SFT. This process is inherently inefficient because it requires generating a massive volume of training data in a single iteration to achieve effective SFT training. According to Dong & Ma (2025b), they generate 75,000 conjectures per iteration, which is larger than the entire statement set for our generation. Conversely, GAR operates within an online RL cycle. This allows the fuser and prover to continuously update more efficiently, achieving superior performance without the computational burden of the large-scale per-iteration data required by STP.
Appendix D Additional experiments
This appendix section provides additional experiments that are omitted in the main paper due to space limitations.
D.1 Results on PutnamBench
Given that Lean4 formal reasoning is a rapidly progressing field, the MiniF2F-Test serves as a benchmark with high accuracy in current models. To further validate that the GAR training can also make the model perform better on more advanced benchmarks, we evaluate both DeepSeek-Prover-V2-7B and GAR-trained DeepSeek-Prover on PutnamBench under pass@32. The results are demonstrated in Table 5.
We can see that the GAR-trained model solves four additional problems compared to the base model on this challenging benchmark. This consistent improvement across MiniF2F, PutnamBench, and ProofNet demonstrates the robustness of the implicit curriculum established by GAR.
D.2 Efficiency Study
| Model Type | PutnamBench |
| DeepSeek-Prover-V2-7B (Ren et al., 2025) | 22/660 |
| GAR DeepSeek-Prover | 24/660 |
| Model Type | Base Model | GAR trained |
| DeepSeek-Prover-V2 (Ren et al., 2025) | 6.96% | 13.11% |
| Goedel-Prover-V2 (Lin et al., 2025b) | 24.18% | 27.05% |
We analyze the computational efficiency of the GAR framework, as training from scratch demands substantial resources, to evaluate this, we compare our approach with Kimina-Prover (Wang et al., 2025a), selected for its transparent reporting of training details. The number of roll-out theorems and proofs per iteration in Wang et al. (2025a) matches our configuration. From Figure 4 in their work, we observe that Kimina-Prover achieves approximately 2% of performance gains after 25 training iterations, where GAR trained Goedel-Prover-V2-8B reaches more improvements in merely three iterations.
Notably, despite Kimina-Prover starting from an SFT base model, while GAR starts from a heavily RL-trained base model, this comparison strengthens our efficiency claim. Because it is well established in prior work (Guo et al., 2025) that models already heavily optimized via RL face diminishing returns and are inherently harder to improve than SFT baselines. The fact that GAR achieves relatively significant gains on top of RL-optimized base models in only 3-5 iterations, while Kimina-Prover gains approximately 2% from an SFT starting point over 25 iterations, demonstrates the high sample efficiency of our adversarial training paradigm. Given resource constraints that prevent exhaustive experimentation, we leave the discussion of the scalability of GAR for future work.
D.3 Statement Modification Rate Study
To dive deeper into the behavior of statement modification, we provide detailed statement modification rates of both the base model and the GAR trained model in MiniF2F-Test under pass@32. The results are shown in Table 6
From the comparison of base models, we can see that the Goedel’s increased statement modification rate happens together with its performance enhancement. We attribute this to the stronger self-correction capability it obtains. Similarly, the soft penalty in GAR ensures that the model is penalized if it simplifies the problem to a triviality. Furthermore, we can observe that when the modification rate is low, the GAR training introduces a higher modification rate, as is the case for DeepSeek-Prover. However, if the modification rate is high in the base model, the soft penalty will effectively control it within a reasonable range. These findings prove the improvement of the GAR-trained model from another point of view
D.4 Additional Ablation Studies
D.4.1 Frozen Fuser Study
To further analyze our generative adversarial training, we conduct this experiment that applies the GAR training only on the prover and keeps the statement fuser untrained. We run three iterations of Frozen fuser GAR training on Goedel-Prover-V2-8B for three iterations. The MiniF2F-Test pass@32 results are demonstrated in Table 7.
From the results, we can see that frozen fuser fails to obtain any performance improvements compared to the base model. This confirms that a static generator is unable to extend the prover’s capabilities beyond its initial limits. This proves the necessity of co-evolution for both the problem composer and the prover.
D.4.2 Single Problem Enhancement Fuiser
In order to further study the effectiveness of our idea of problem fusing, which enhances the problem difficulty by combining two problems into a single one. We conduct the experiment of replacing the statement fuser with MagicCoder (Wei et al., 2023) that enhances the problem difficulty by restating the provided single problem to make it more advanced. The results are presented in Table 7.
The result indicates that replacing the fusion process with the MagicCoder style problem-evolving strategy results in a 2.46% performance drop compared to the base model. We conclude this degradation to the limited capability of Qwen3-8B to create more advanced problems with only one reference. It may not provide sufficient difficulty gradient for the prover, leading to training failure. In contrast, our fusion approach creates a more effective implicit curriculum by combining distinct concepts from multiple problems, thereby validating the design of the Statement Fuser.
| Method | MiniF2F-Test |
| Base model | 77.87% |
| Frozen Fuser | 77.87% |
| MagicCoder Fuser | 75.41% |
| GAR trained | 80.33% |
D.5 Case study
D.5.1 Case 1:
In this section, we demonstrate an example of how the GAR trained model successfully solves the problem while the base models fail. Compared to the base prover, the GAR trained model relies less on brute simplification and more on early structural framing, producing shorter, steadier proof traces that remain faithful to the statement and are more likely to pass the verification. If we dig deeper into the Long CoT of two models, we can find that GAR-trained prover writes a shorter and more focused long CoT. While the base model’s CoT is longer but less directed, it cycles among general identities and simplifications without stabilizing on a closing invariant. The result that GAR’s trace is both and clearer and more verifiable, which leads to the successful proof.
GAR-trained model generated result:
Base model generated result:
D.5.2 Case 2:
In another ProofNet example, GAR again succeeds where the base prover fails. This instance more clearly illustrates how GAR mitigates useless simplifications. It uses NLs to reach a comprehensive proof plan and effectively conduct the proof in a relatively simple manner. However, the base model conducts a quicker search but a prolonged proof. It expends effort on low-yield algebra and yields lengthy, fragile derivations. These observations highlight that the implicit curriculum in GAR fosters more efficient reasoning paths and greater verification robustness.
GAR-trained model generated result:
Base model generated result: